OpenStudy (cderaalien):

help Josh and his friends bought vanilla wafers for $4 per packet and chocolate wafers for $1 per packet at a carnival. They spent a total of $45 to buy a total of 27 packets of wafers of the two varieties. Part A: Write a system of equations that can be solved to find the number of packets of vanilla wafers and the number of packets of chocolate wafers that Josh and his friends bought at the carnival. Define the variables used in the equations. Part B: How many packets of chocolate wafers and vanilla wafers did they buy? Explain how you got the answer and why you selected a particular

OpenStudy (texaschic101):

v + c = 27 4v + c = 45 these are your 2 equations...you can easily solve by substitution or elimination

OpenStudy (anonymous):

aaaand he beat me to it :/

OpenStudy (cderaalien):

Yes, but how did you get those equations exactly? That's the part I always have trouble on.

OpenStudy (anonymous):

so to get the 45$ you will have an amount of each wafer (variables) and the price corresponding with it will get you to the actual dollar value.

OpenStudy (texaschic101):

first equation we are just dealing with the total number of packets. v + c = 27.....vanilla + chocolate = 27 second equation is dealing with the cost of the packets.. 4v + c = 45......$4 per pack of vanilla and $1 per pack of chocolate = $45

OpenStudy (anonymous):

so if we look at it the reverse way, you have 10 vanilla wafers and 4 chocolate. 4(10) + 1(4) = 44 (this isnt meant to answer your actual problem)

OpenStudy (cderaalien):

Thank you @texaschic101

OpenStudy (cderaalien):

Thank you also @whalen

OpenStudy (texaschic101):

glad to help :)

OpenStudy (anonymous):

np, if only texaschic wasnt a bit faster than me i wouldve been more of a help lol

OpenStudy (cderaalien):

but wait, what does the variables stand for?

OpenStudy (cderaalien):

does v = vanilla and c = chocolate?

OpenStudy (cderaalien):

@pooja195 does the variables mean what I thought?

pooja195 (pooja195):

Yes

OpenStudy (cderaalien):

Okay, thank you :). Just wanting to make sure they didn't mean the total of them or something.

pooja195 (pooja195):

I see yw ^_^